# Shahriar Noroozizadeh

### v1

@@ Abstract
In sequence modeling, the parametric memory of atomic facts has been predominantly abstracted as a brute-force lookup of co-occurrences between entities. We contrast this associative view against a geometric view of how memory is stored. We begin by isolating a clean and analyzable instance of Transformer reasoning that is incompatible with memory as strictly a storage of the local co-occurrences specified during training. Instead, the model must have somehow synthesized its own geometry of atomic facts, encoding global relationships between all entities, including non-co-occurring ones. This in turn has simplified a hard reasoning task involving an $\ell$-fold composition into an easy-to-learn $1$-step geometric task.
From this phenomenon, we extract fundamental aspects of neural embedding geometries that are hard to explain. We argue that the rise of such a geometry, despite optimizing over mere local associations, cannot be straightforwardly attributed to typical architectural or optimizational pressures. Counterintuitively, an elegant geometry is learned even when it is not more succinct than a brute-force lookup of associations.
Then, by analyzing a connection to Node2Vec, we demonstrate how the geometry stems from a spectral bias that—in contrast to prevailing theories—indeed arises naturally despite the lack of various pressures. This analysis also points to practitioners a visible headroom to make Transformer memory more strongly geometric. We hope the geometric view of parametric memory encourages revisiting the default intuitions that guide researchers in areas like knowledge acquisition, capacity, discovery and unlearning.
@@ 1 Introduction
Neat representations materialize when a deep network needs to compress redundancies in data. On the other extreme, if the data is a set of atomic facts (like the birth date of a celebrity), the network would simply memorize these incompressible associations as a lookup table [126, 17]. These two narratives have so far roughly guided our understanding of how neural networks fit sequential data. This paper fleshes out a third phenomenon in sequence modeling—glimpses of which were observed recently in Khona et al. [76], Ye et al. [172]—where a neat representation materializes from memorizing incompressible atomic facts. We argue that this phenomenon implies a geometric form of parametric memory, one that encodes global relationships between non-co-occurring entities. We point out that this is dramatically different from the common associative view of parametric memory as storing mere local atomic co-occurrences. From this geometry of atomic facts, we extract aspects of neural geometries that are hard to explain, posing fundamental questions about memorization in deep sequence models. To these questions, we offer some preliminary answers.
@@
Our discussion starts at a seemingly tangential point. We begin by consolidating a fragmented set of recent demonstrations [76, 43, 172, 158] that the Transformer exhibits some level of implicit in-weights reasoning, i.e., reasoning over knowledge from the weights without emitting an explicit chain of thought. We sharpen these results by crafting a scenario where this ability is less expected, plays out vividly, and can be cleanly isolated and analyzed. Specifically, we study path-finding on path-star graphs, a (symbolic) implicit reasoning task. The task was adversarially designed [12] to cause failure of next-token trained deep sequence models—both the Transformer [156] and Mamba [54] alike. Whereas in the original task, the model is given the graph in-context, here we make the model memorize the graph’s edges in its weights. Whereas in the original task, the models spectacularly fail to learn path-finding even on small graphs, in our in-weights task, our models (both the Transformer and Mamba) succeed even on massive graphs.
@@
Figure 1: Associative vs. geometric memory of models trained on various graphs. Parametric memory in deep sequence models is often abstracted as if co-occurrences from atomic facts are stored in a weight matrix, while the co-occurring entities themselves are embedded arbitrarily [161, 46, 70, 183, 23, 111, 22, 17, 160]. (left). This associative view is hard to reconcile with our observation that the Transformer learns implicit reasoning on an in-weights path-star graph. In practice, the learned embeddings (middle) reflect global structure inferred from the local co-occurrences, implying a geometric view of memory. Both these views are valid ways to fit the training data, but it is not straightforward why the latter prevails during Transformer optimization. When associative memory is explicitly prohibited, as in a Node2Vec model (right), a more elegant geometry materializes. This points to a clear headroom to improve the geometric nature of a Transformer’s memory. Details of the Transformer architecture used for this visualization are provided in §B.2.2. Similar geometries for Mamba and neural networks are presented in §C.3.
@@
The success in the in-weights path-star task, we argue, is hard to reconcile within the associative view of parametric memory, a convenient and highly effective abstraction of neural network memory, popular in literature [17, 22, 70, 161, 46, 183, 23, 111, 160]. In this abstraction, knowledge is stored in a matrix, say ${\bf{W}}_{\tt assoc}$ such that if $u$ and $v$ co-occur in the same context during training, the quantity ${\bf{\Phi}}(v)^{T}{\bf{W}}_{\tt assoc}{\bf{\Phi}}(u)$ is large, for some arbitrary embedding $\bf{\Phi}$}. With such a data structure, the path-star task requires composing the aforementioned matrix operation $\ell$ times (for path length $\ell$). Unless there is step-wise supervision for each composition, learning the $\ell$-fold composition should intuitively be a daunting needle-in-the-haystack task demanding $\exp(\ell)$ time. By the specific design of our task, every possible form of step-wise guidance is eliminated. Yet our model appears to find the needle.
@@
This apparent paradox begins to be resolved due to salient observations in Khona et al. [76], Ye et al. [172] that the Transformer represents the nodes of their graphs with some notion of distance. We generalize this observation, pointing out that this has profound implications and is in fact non-trivial to explain. First, the existence of such node embeddings implies that atomic facts can be stored in an altogether distinct paradigm of parametric memory, a geometric one, which is in dramatic contrast to the associative one. In the simplest form of this view, co-occurrences are not simply stored in a matrix $\bf{W}$}, but rather neatly encoded within the embeddings, say ${\bf{\Phi}}_{\tt geom}$ (see Fig.˜1). In fact, these embeddings are so carefully arranged that they encode global relationships without explicit supervision to do so: even if entities $u$ and $v$ never appeared in the same context, the dot-product ${\bf{\Phi}}_{\tt geom}(u)^{T}{\bf{\Phi}}_{\tt geom}(v)$ captures the model’s own notion of multi-hop distance between the two. This has implications on reasoning since what seemed to be a hard-to-learn $\ell$-fold composition of local associations, now becomes an easy-to-learn $1$-step geometric task for the model.
@@
While helping make sense of our paradox, the observed geometry raises fundamental questions. First, there must be competition between the two parametric memories, both equally valid solutions to the training objective; why does the geometric prevail over the associative? To some readers, a geometric bias may seem familiar and intuitive at first sight; but we isolate aspects that cannot be easily explained. Counterintuitively, we argue that an elegant geometry is not necessarily more succinct than a lookup of local associations. Thus, typical capacity pressures from the architecture or optimization do not explain why a highly non-trivial geometry is synthesized—that too from optimizing only over local associations. Towards understanding this, we draw connections to the simpler Node2Vec architecture, where we find that global geometries emerge from well-known spectral biases in such architectures. But, in contrast to prevailing theories (e.g., Levy and Goldberg [84]; see §5.6), we identify how the spectral bias arises naturally, independent of typically-assumed architectural or optimizational pressures. This gives us a preliminary insight into the natural rise of a geometric memory, albeit in a much simpler model, leaving open a foundational question for deep sequence models.
1.1 Implications.
@@
Although the evidence of implicit reasoning is so far limited to symbolic tasks, we believe that the geometry we isolate is a clean nucleus of geometries known to arise in language modeling e.g., [36, 102, 58]. The insights from this nucleus helps conceive broader directions, practical and theoretical. First, we find that the embeddings learned by the more naive Node2Vec models are more strongly geometric than that of Transformers; in hindsight, this points to a well-specified headroom for making Transformer memory more geometric and less associative in practice. If the geometric bias can be improved in natural language tasks, it would also benefit natural implicit reasoning tasks where results have so far been mixed [122, 18, 169, 170, 14]. Since geometric memory encodes global relationships, it could pave the way to combinational creativity [19, 40, 108]: discovering novel connections between information scattered in a large pretraining set. On the flip side, the interdependencies in geometric storage may impose limits on knowledge editing, unlearning and accurate retrieval. Another implication of our study is a support for why parametric memory may be superior to in-context memory, echoing Wang et al. [158], Geerts et al. [43]. Finally, the gap between the Transformer and Node2Vec geometries may also be of interest to practitioners in retrieval systems, when choosing between modern generative retrieval models [150, 163, 128] and traditional dual-encoder models [47, 66].
@@
Many foundational directions also follow. Foremost are the questions of how associative and geometric memory compete under gradient descent and what optimization settings are ideal for the geometry to arise. The examples and analyses in this work may act as sandboxes to study a variety of geometric empirical phenomena: the emergence of “world models” [39, 58, 138], linear representations and superposition in interpretability [102, 117, 36] and why different models share similar representations, i.e., the Platonic representation hypothesis [83, 67]—and what these representations are. Finally, we speculate that the associative view forms the default unstated set of intuitions that guide research in numerous areas such as knowledge acquisition, discovery, unlearning, reasoning and storage capacity; the geometric view may inspire researchers to revisit, spell out and widen these latent intuitions.
@@ 1.2 Summary of contributions
    We isolate a clean, analyzable instance of implicit in-weights reasoning, and contrast the predominant local associative memory against a global geometric view of memory in deep sequence models (simple neural networks, Transformers and Mamba).
@@
    We show why this observation is surprising, arguing that the emergence of the geometric memory over the associative memory cannot be attributed to obvious architectural, optimizational or supervisory pressures.
@@
    We connect the global geometry to the spectral bias of Node2Vec dynamics and empirically intuit how it emerges without typically assumed pressures. This makes progress towards an open question in Node2Vec and highlights significant headroom in the embedding geometry of the current architectures.
2 Experiments: Implicit in-weights reasoning is learned
@@
We investigate planning on a path-star graph, a task designed to be adversarial towards next-token learning [12]. The task has a clear notion of a chain-of-thought, and a well-understood mechanism of failure for learning in-context reasoning; our hope is to repurpose this to cleanly analyze in-weights reasoning in a way that was not possible in earlier studies.
@@
Background. The path-star topology (Fig.˜2) consists of a root node with multiple disjoint (uniform-length) paths branching outwards. In the version of B&N’24, a model is given in context the adjacency list (with randomized node labeling and edges ordering, but fixed topology) and a goal leaf node; the task is to predict the unique path from the root to a specified goal node, without explicitly producing a chain-of-thought. To succeed, one merely needs to notice a simple right-to-left structure: the solution is the unique path back from the leaf node, reversed. Indeed, this is the implicit chain-of-thought required before emitting the first token.
@@
Yet, on this simple task, left-to-right next-token learners are known to fail in-distribution. This failure unfolds in two stages during training: (1) On all but the first token, the model learns a trivial solution (termed a Clever Hans cheat) that is much simpler than planning: simply predict the token as the unique child of the previous ground-truth token that is revealed in the context. This crucially starves the first, key decision-making step of gradients from the rest of the path. (2) Consequently, the first token must be learned in isolation, which becomes a computationally hard learning $\ell$-hop composition problem (as detailed later). At test-time, this model defaults to guessing a random first token and continuing along that wrong path. More details of this failure are in §A.
2.1 In-weights path-star task.

### v2

@@ Abstract
Deep sequence models are said to store atomic facts predominantly in the form of associative memory: a brute-force lookup of co-occurring entities. We identify a dramatically different form of storage of atomic facts that we term as geometric memory. Here, the model has synthesized embeddings encoding novel global relationships between all entities, including ones that do not co-occur in training. Such storage is powerful: for instance, we show how it transforms a hard reasoning task involving an $\ell$-fold composition into an easy-to-learn $1$-step navigation task.
From this phenomenon, we extract fundamental aspects of neural embedding geometries that are hard to explain. We argue that the rise of such a geometry, as against a lookup of local associations, cannot be straightforwardly attributed to typical supervisory, architectural, or optimizational pressures. Counterintuitively, a geometry is learned even when it is more complex than the brute-force lookup.
Then, by analyzing a connection to Node2Vec, we demonstrate how the geometry stems from a spectral bias that—in contrast to prevailing theories—indeed arises naturally despite the lack of various pressures. This analysis also points out to practitioners a visible headroom to make Transformer memory more strongly geometric. We hope the geometric view of parametric memory encourages revisiting the default intuitions that guide researchers in areas like knowledge acquisition, capacity, discovery, and unlearning.
@@ 1 Introduction
When succinct high-level patterns explain the data, deep sequence models produce corresponding high-level representations, as seen in natural language [40, 130, 112] and arithmetic tasks [100, 120, 117, 59]. When no such patterns exist (e.g., as in the capitals of countries), models are said to default to a brute-force lookup, known as associative memory [139, 20]. These two narratives have so far roughly guided our understanding of how neural networks fit sequential data. We highlight a third behavior missing in this narrative: even when memorizing such incompressible atomic co-occurrences, a deep sequence model can produce powerful representations. This implies a distinct form of parametric memory that we term as geometric memory—glimpses of which can be found in recent observations [74, 85, 189]. In opposition to the well-studied associative memory, a geometric memory has synthesized “global” information not explicit in the local co-occurrences, enabling new forms of reasoning. From this behavior, we extract aspects of neural embedding geometries that are hard to explain, raising fundamental questions about memorization in deep sequence models. To these questions, we offer some preliminary answers.
@@
Our discussion starts at a seemingly tangential point. We begin by consolidating a fragmented set of recent demonstrations [85, 174, 43, 48, 189, 74] that the Transformer exhibits some level of implicit in-weights reasoning, i.e., reasoning over knowledge from the weights without emitting an explicit chain of thought. We sharpen these results by crafting a scenario where this ability is unexpected, plays out vividly, and can be cleanly isolated and analyzed. Specifically, we study path-finding on path-star graphs, a (symbolic) implicit reasoning task. The task was adversarially designed [13] to cause failure of next-token trained deep sequence models—Transformer [172] and Mamba [60] models alike. Whereas in the original task, the model is given the graph in-context, here we make the model memorize the graph’s edges in its weights. Where before the model spectacularly failed to learn path-finding even on small graphs, in our in-weights task the model succeeds even on massive graphs.
@@
Figure 1: Associative vs. geometric memory of models trained on various graphs. There are two dramatically different ways to memorize a dataset of atomic facts. The common view is of associative memory: entities are embedded arbitrarily, and co-occurrences are stored in weight matrices. (left). §2.4: In practice, we find a geometric memory: the learned embeddings of a Transformer (middle) reflect global structure inferred from the local co-occurrences in training data. §4: When associative memory is explicitly prohibited (by removing intermediate layers), as in a Node2Vec model (right), a more elegant geometry materializes. This points to a clear headroom to improve the geometric nature of a Transformer’s memory. Details of the Transformer architecture used for this visualization are provided in §B.2.2. Similar geometries for Mamba and neural networks are presented in §C.3.
@@
The success in the in-weights path-star task, we argue, is hard to reconcile within the associative view of parametric memory, a convenient and highly effective abstraction of neural network memory, popular in literature [20, 26, 78, 177, 51, 202, 27, 124, 176]. In this abstraction (see Section˜2.3), each entity is embedded randomly/near-orthogonally via a “key” function ${\bf{\Phi}}(\cdot)$, while associations are stored in what is effectively a (rotated) adjacency matrix, ${\bf{W}}_{\tt assoc}$. To perform a “lookup”, one computes the quantity ${\bf{\Phi}}(v)^{T}{\bf{W}}_{\tt assoc}{\bf{\Phi}}(u)$ which is large if and only if entities $u$ and $v$ co-occur during training. With such a data structure, our path-finding task requires composing the lookup operation $\ell$ times (for path length $\ell$). Unless there is step-wise supervision for each composition, learning the $\ell$-fold composition should be a daunting needle-in-the-haystack task demanding $\exp(\ell)$ time. By the specific design of our task, every possible form of step-wise guidance is eliminated. Yet our model appears to find the needle.
@@
This apparent paradox begins to be resolved by a preliminary observation in [78, 85, 189]: that the node embeddings $\bf{\Phi}$} reflect some useful notion of distance; and a separate finding in [74, 153]: that (two-hop) reasoning can emerge from “factorized” parameterizations. We flesh out a unified understanding of these observations and their implications. First, the existence of structured node embeddings implies that models can store atomic facts in an altogether distinct paradigm of parametric memory that we term as geometric memory. This memory is in dramatic contrast to the associative one. In the simplest form of the geometric view (see Section˜2.4), the associations are no longer stored verbatim in a lookup table ${\bf{W}}_{\tt assoc}$, but are (low-rank) factorized into the embeddings as ${\bf{\Phi}}_{\tt geom}(u)^{T}{\bf{\Phi}}_{\tt geom}(v)$. Our key insight is that these embeddings are no longer arbitrary. Rather, they are carefully arranged such that they encode global relationships without explicit supervision to do so: even if entities $u$ and $v$ never appeared in the same context, the dot-product ${\bf{\Phi}}_{\tt geom}(u)^{T}{\bf{\Phi}}_{\tt geom}(v)$ captures the model’s own notion of multi-hop distance between the pair. This global geometry can open up new forms of reasoning, since what seemed to be a hard-to-learn $\ell$-fold composition of local associations is now approximated as an easy-to-learn spatial navigation task. We contrast the associative (Section˜2.3) and geometric (Section˜2.4) views of memory visually in Fig.˜1 and also in Table˜1.
@@
The observed geometry raises fundamental questions. Most importantly, there must be competition between the two parametric memories, both equally valid solutions to the training objective; why does the geometric prevail over the associative? To some readers, a geometric bias may seem familiar and intuitive at first sight, but we isolate aspects that cannot be easily explained: the global geometry is learned even when there is no “global" supervision (e.g., a path-finding objective), even when there is no rank constraint, even when the geometry is much harder to find, and even when it is just as succinct as a lookup. Thus, typical pressures from the supervision, architecture or optimization do not explain why a highly non-trivial geometry is synthesized. We term this the memorization puzzle. Towards understanding this, we reduce the setup to a minimal Node2Vec architecture, where we find that global geometries emerge from well-known spectral biases in such architectures. But, in contrast to prevailing theories (e.g., Levy and Goldberg [93]; see §5.5), we identify how the spectral bias arises naturally from cross-entropy loss minimization, independent of typically-assumed architectural or optimizational pressures. This gives us a preliminary insight into the natural rise of a geometric memory, albeit in a much simpler model, leaving open a foundational question for deep sequence models. Various practical questions open up too, which we discuss below.
1.1 Implications
@@
Although the evidence of implicit reasoning is so far limited to symbolic tasks, we believe that the geometry we isolate is a clean nucleus of well-established geometries in language modeling [40, 112, 65] and arithmetic tasks [100, 120, 117, 59]. These insights suggest broader directions, both practical and theoretical. First, we find that the embeddings learned by more naive Node2Vec models are more strongly geometric than those of Transformers; in hindsight, this points to a well-specified headroom for making Transformer memory more geometric and less associative in practice. If the geometric bias can be improved in natural language tasks, it would also benefit natural implicit reasoning tasks where results have so far been mixed [135, 21, 185, 186, 16]. Since geometric memory encodes global relationships, it could help discover novel connections between information scattered in a large pretraining set [187] and thus, some types of creativity [22, 45, 119]. On the flip side, the interdependencies in geometric storage may impose limits on knowledge editing, unlearning and accurate retrieval; it may also hallucinate associations [74] (cf. “illusory correlations” in [153]). Another implication of our study is a support for why parametric memory may be superior to in-context memory, echoing Wang et al. [174], Geerts et al. [48]. Finally, the gap between the Transformer and Node2Vec geometries may also be of interest to practitioners in retrieval systems, when choosing between modern generative retrieval models [166, 179, 141] and traditional dual-encoder models [52, 73].
@@
Many foundational directions follow. Foremost are the questions of how associative and geometric memory compete under gradient descent and what aspects of the data or optimization are ideal for the geometry to arise. The examples and analyses in this work may act as sandboxes to study a variety of geometric empirical phenomena: the emergence of “world models” [44, 65, 151], linear representations and superposition in interpretability [112, 130, 40] and why different models share similar representations, i.e., the Platonic representation hypothesis [92, 75]—and what these representations are. Finally, we speculate that the associative view forms the default unstated set of intuitions that guide research in numerous areas such as knowledge acquisition, discovery, unlearning, reasoning, storage capacity and scaling laws. The geometric view may inspire researchers to revisit, spell out and widen these latent intuitions. For instance, one may re-examine capacity [106, 81, 105, 80, 87] and scaling laws [10, 115, 144, 128] under a geometric lens and find very different answers.
@@ 1.2 Summary of contributions
    We isolate a clean instance of implicit in-weights reasoning in deep sequence models that succeeds even without any form of stepwise supervision. This isolates a behavior that is hard to explain within the predominant associative view of parametric memory. (§2)
@@
    We establish an alternative, global geometric view of memory in deep sequence models (simple neural networks, Transformers and Mamba). (§2.4)
@@
    We demonstrate why the geometric memory is surprising: the emergence of geometric over associative memory cannot be attributed to obvious architectural, optimizational or supervisory pressures. With this, we formulate a memorization puzzle in sequence modeling. (§3)
  4. 4.
    We connect the global geometry to the spectral bias of Node2Vec dynamics. We empirically intuit how it emerges without typically-assumed pressures, when minimizing the cross-entropy loss. This makes progress towards an open question in Node2Vec. We also highlight significant headroom in the embedding geometry of the current architectures. (§4)
Table 1: Comparison of two forms of parametric memory in a deep sequence model, given a dataset of edges from a graph $\mathcal{G}=({V},E)$
Associative Memory (Section˜2.3) Geometric Memory (Section˜2.4)
Storage mechanism A lookup table of pairwise edge associations A graph embedding of the vertices
Embedding $\bf{\Phi}$} Arbitrary (random or orthogonal); acts as keys for lookup; doesn’t reveal the associations in itself Reveals the graph structure
Simplest neural implementation (of the learned similarity between input token, node $u$ and next-token, node $v$, denoted by $f(u)[v]$) $f(u)[v]={\bf{\Phi}}(u)^{T}{\bf{W}}_{\tt assoc}{\bf{\Phi}}(v)$ where ${\bf{W}}_{\tt assoc}\approx\underset{(u,v)\in\mathcal{G}}{\sum}{\bf{\Phi}}(u)\otimes{\bf{\Phi}}(v)$} and ${\bf{\Phi}}(u)\cdot{\bf{\Phi}}(v)\approx 0$ for $u\neq v$. $f(u)[v]={\bf{\Phi}}_{\tt geom}(u)^{T}{\bf{\Phi}}_{\tt geom}(v)$ where ${\bf{\Phi}}_{\tt geom}$ corresponds to (top) eigenvectors of graph Laplacian.
Relationship to graph $\mathcal{G}$ ${\bf{W}}_{\tt assoc}$ can be thought of as adjacency matrix (up to a rotation) ${\bf{\Phi}}_{\tt geom}$ can be thought of as low-rank factorization of adjacency matrix; alternatively, the eigenvectors in ${\bf{\Phi}}_{\tt geom}$ are also components dominant in multi-hop adjacency matrix, $\bf{A}$$^{\ell}$ for large $\ell$.
Global information Only local edge information is readily available i.e., $f(u)[v]$ high only for $(u,v)\in E$ Global information readily available i.e., $f(u)[v]$ is proportional to multi-hop distance between $u$ and $v$ in $\mathcal{G}$
Novel information not present in data None; only stores seen associations. Has synthesized novel information by integrating information scattered across multiple datapoints.
Implicit compositional reasoning e.g., path-finding Intuitively, exponentially hard to learn. Approximated as an easy-to-learn geometric navigation task.
Accuracy of retrieval Precise Approximate
Description length Number of edges, $|E|$ Number of vertices times the embedding dimensionality, |V|m|{V}|m
2 Implicit in-weights reasoning is learned
@@
We investigate planning on a path-star graph, a graph designed to be adversarial towards next-token learning [13]. The task has a clear notion of a chain-of-thought, and a well-understood mechanism of failure for learning in-context reasoning. Thus, we hope to repurpose this to cleanly analyze a different type of reasoning—in-weights reasoning—in a way that was not possible in earlier studies of in-weights reasoning.
@@
Background. The path-star topology (Fig.˜2) from B&N’24 is a tree graph where multiple disjoint uniform-length paths branch outwards from the root. In other words, all internal nodes but the root have exactly one child. In the original version of this task [B&N’24], the model is given a fresh example of this graph in context as an adjacency list (with randomized node labeling and edge ordering, but fixed topology), along with a leaf node specified as goal. The model must spell out the nodes in the unique path from the root to the goal in that order. The solution is to notice a simple right-to-left structure: take the unique path back from the leaf node and reverse it. Indeed, this is the implicit chain-of-thought required before emitting the first token.
@@
Yet, on this simple task, left-to-right next-token learners are known to fail in-distribution even on tiny graphs. This failure unfolds in two stages during optimization: (1) On all but the first token, the model fits the data trivially: simply predict the token as the unique child of the previous ground-truth token that is revealed in the context. Then, this starves the first, key decision-making step of gradients from the rest of the path. (2) Consequently, the first token must be learned in isolation, which becomes a computationally hard learning $\ell$-hop composition problem (as detailed later). At test-time, this model defaults to guessing a random first token and continuing along that wrong path. We reproduce this failure in Fig.˜5 and elaborate on it in §A.
2.1 The in-weights path-star task